|
AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
|
CARME |
 |
|
Message ID: 375
Entry time: Fri Aug 12 09:06:57 2022
|
Author: |
TD, JM |
Subject: |
Noise tests pt2 |
|
|
10.07 DSSSD bias -100V leakage current -8.65uA
Waveforms and discriminators off, still on clean power supply
stats are ~10k across all fees
aida03 ~ 220
aida04 ~ 205
aida01 ~ 378
aida02 ~ 259
aida05 ~ 40
Grounding removed from all fees except nn bias fee aida04 and aida05. Grounding goes into both adaptor board and fee body
aida01 - no identifiable peak
aida02 - no identifiable peak
aida05 - double peak separated by 70 chn
aida04 - 400 shouldering
aida03 - no identifiable peak
Grounding removed from all fees except nn and pn bias fees aida04, aida01 and aida05. Grounding goes into both adaptor board and fee body
aida01 - 405
aida02 - no identifiable peak
aida05 - double peak
aida04 - 228
aida03 - 417
Grounding removed from all fees except pn bias fee aida01 and aida05. Grounding goes into both adaptor board and fee body
aida01 - 383
aida02 - no identifiable peak
aida05 - 64
aida04 - no identifiable peak
aida03 - no identifiable peak
Grounding on all fees. Grounding goes into adaptor board only
aida01 - no identifiable peak
aida02 - no identifiable peak
aida05 - 63 (wasn't changed)
aida04 - 700
aida03 - 700
Back to previous grounding on all fees
aida01 - 331
aida02 - 226
aida05 - 37
aida04 - 207
aida03 - 214
power cycle
aida01 - 326
aida02 - 238
aida05 - 46
aida04 - 212
aida03 - 195
Waveforms and discriminator back on
Removed test inputs and terminators from all fees, except aida03 (unable to reach) however the lemo is not touching the chamber
See attached waveforms compared to previous elog 32,33
Inserted test input to aida05 only directly from pulser. Width ~ 28
Inserted test input to aida01 only directly from pulser. Width ~ 329
Waveforms and discriminator off -> no impact on aida01 pulser width
Waveforms and discriminator back on
Connected a grounding cable from one of the water cooling pipes to cryring ground
aida01 ~ 205
aida05 ~ 23
aida02 ~ 105
aida03 ~ 71
aida04 ~ 148
Fees with bias attached to them aida01 (+ve) and aida04 (-ve) both show larger widths than there polarity counterpart
waveform screencap 34,35
Fee probe
7.46V left most bar
-6.26V 2nd bar
5.26V 3rd bar
5.26V 4th bar
5.25V 5th bar
Power supply probe
Top -6.05
2nd -6.05
3rd -6.05
4th +6.49
5th -7.78
new Ortec 448 pulser swapped in
aida04 ~ 166
aida05 ~ 21
Attachment 7-11 oscilloscope picture when probing fee relay power supply. Power ON fees connected. Top is attachment 7, bottom (5th) is attachment 11.
Attachment 12-16 old pulser
Attachment 17-18 new pulser
Attachment 19-23 oscilloscope picture when probing fee relay power supply. Power OFF fees connected. Top is attachment 19, bottom (5th) is attachment 23.
Attachment 24-28 oscilloscope picture when probing fee relay power supply. Power ON, fees disconnected. Top is attachment 24, bottom (5th) is attachment 28.
Attachment 29,30 probe-scope setup |
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|